You are here: Foswiki>LabstaffWeb Web>LabstaffDowntime>Outage20100718MRI (20 Jul 2010, lindss2@LAB.CS.RPI.EDU)EditAttach

MRI cooling outage


  • OPS paged at 11:20pm on the 17th saying the cooling had failed and the temperature was going up.
  • No word from HVAC as of midnight, so cluster shut down
  • Machines restored to service 7/19/2010
Topic revision: r2 - 20 Jul 2010, lindss2@LAB.CS.RPI.EDU
 

This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback